OpenStudy (anonymous):

A computer chooses a point at random from either the blue or the red triangles. The theoretical probability of a point being chosen from the red triangle is _____. After 8 trials, the computer has chosen a point inside the red triangle 5 times. The experimental probability of choosing a point from the red triangle is _____. Fill in the blanks in the previous sentences to make the statements true. Then use your answers to help answer the following question. Should you doubt whether the computer's choices are actually random?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it would help if we had a picture of the triangles are they of equal areas or different?

OpenStudy (misty1212):

looks like half the area is red, so the probability of picking red is one half i.e. \(\frac{1}{2}\)

OpenStudy (anonymous):

that would be the experimental probability? it says it could be greater than .5 or between .2 and .5

OpenStudy (misty1212):

The theoretical probability of a point being chosen from the red triangle is _____. one half

OpenStudy (misty1212):

After 8 trials, the computer has chosen a point inside the red triangle 5 times. The experimental probability of choosing a point from the red triangle is _____. \[\frac{5}{8}>\frac{1}{2}\]

OpenStudy (anonymous):

ohh I understand, thanks for your help I have been stuck forever.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]